# Ultra-wideband Time Difference of Arrival Indoor Localization: From Sensor Placement to System Evaluation

### v1

@@ IV-A Multi-agent Pedestrian Localization
We first demonstrate the accuracy and robustness of our UWB TDOA positioning system in multi-agent pedestrian localization through comprehensive real-world experiments. We used eight DWM1000 UWB radios as anchors to set up the UWB TDOA localization system. The anchors were configured to be in centralized TDOA mode and communicated in a round-robin network topology with $\Gamma=\{(8,1),(1,2),\cdots,(7,8)\}$. A detailed explanation of centralized and decentralized TDOA modes using Bitcraze’s UWB anchors can be found in [20]. Given that the experimental space is an open, two-story indoor flight arena, we put the anchors at the corners of the space to maximize the coverage of the positioning area (see Figure 3b). The standard deviation of the UWB TDOA measurement noise is set to be $\sigma_{ij}=0.1$ m, which is the default DWM1000 UWB radio precision. We applied a chi-squared test in the ESKF algorithm as the outlier rejection mechanism with the Mahalanobis distance parameter set to $5$. We treat the millimeter-level positioning provided by a Vicon motion capture system as the ground truth. Although the UWB TDOA localization system can provide positioning services to an unlimited number of agents, we demonstrate the multi-agent pedestrian localization performance using five identical handheld devices in our experiments.
@@
We conducted similar experiments using two different anchor constellations. In the first constellation, four anchors were placed on the tripods at a height of $2.65$ meters, leading to a narrower separation in the z-axis. Over three experiment trials, five participants each carried a handheld device and navigated within the convex hull of the anchor constellation. While walking and running freely, the students waved the handheld devices, engaging in a variety of movements within the space. We demonstrate one trial of our experiments together with a comparison of the estimated and the ground truth trajectories for each agent in Figure 4. We invite readers to view our supplementary video (http://tiny.cc/uwb_tdoa_ram24) to gain a better insight into our experimental process and evaluate the robustness of the localization performance. Across the three experimental trials, our localization system demonstrated consistent pedestrian localization performance for the five agents, achieving an average RMSE of $21$ centimeters with a small standard deviation of $1.56$ centimeters. In the second anchor constellation, we mounted the four anchors on the room’s ceiling frame, increasing the z-axis separation to $5.5$ meters. Similarly, three students carried the handheld devices and conducted three trials of experiments by walking around the same room. With identical parameters applied in the ESKF algorithm, the average RMSE for the three experiments in the second constellation is $16$ centimeters with a similar standard deviation of $1.63$ centimeters.
@@
Through these comparative experiments, we observe that changing the sensor placement alone leads to a notable impact on the localization performance, which highlights the importance of UWB anchor placement. We calculate the theoretical lower bounds of localization RMSE in the aforementioned two anchor constellations using sensor placement analysis. The average theoretical RMSE lower bounds in the first and second anchor constellations are $12$ centimeters and $8$ centimeters, respectively. We summarize the average experimental RMSE values for multi-agent pedestrian localization and the theoretical RMSE lower bounds in Figure 5 for comparison. The experimental results agree with the theoretical analysis that sensor placements with lower theoretical RMSE bounds generally yield better localization performance.
@@ IV-B Localization under Sensor Occlusion
One distinctive feature of UWB-Inertial localization is that the sensor set can be partially or completely occluded thanks to the obstacle-penetrating ability of UWB measurements. To demonstrate this feature, we put one UWB handheld device into a backpack (see Figure 6) and conducted three trials of experiments in the aforementioned first anchor constellation. Since the device was completely occluded by the backpack throughout the experiments, we created a motion capture tracking object for the backpack to record the ground truth position data approximately. As before, we summarize the average localization RMSE over three experimental trials together with the theoretical lower bound value in Figure 5. Readers are encouraged to view our supplementary video to gain an intuitive understanding of this experimental process.
The average RMSE from the three experiments is $22$ centimeters, which closely matches the performance of multi-agent pedestrian localization ($21$ centimeters) within the same constellation. Although the fabric backpack did not pose a severe NLOS challenge for UWB measurements, conventional localization algorithms relying on cameras or LiDARs are unsuitable for this application. This result demonstrates that the UWB-inertial localization system has the potential to provide reliable positioning performance even under complete sensor occlusion.
@@ IV-C Localization in a Staircase Environment
Realistic working conditions, such as building monitoring or multi-room scenarios, often pose geometric challenges that can impact UWB TDOA localization performance. The presence of challenging anchor-tag geometry in such environments often leads to degradation in localization accuracy. Consequently, evaluating localization performance through sensor placement analysis is crucial to avoid areas with poor positioning accuracy. To showcase this ability, we conducted experiments in a geometrically challenging staircase environment. We illustrate how sensor placement analysis effectively captures the degradation of localization accuracy induced by the difficult anchor-tag geometry in staircase positioning scenarios.
Refer to caption (a) (b) (c)
Figure 7: The staircase scenario and our experimental setup are shown in (a) and (b). The experimental trajectory #$\#1$ and the corresponding ground truth trajectory are shown in (a). The ESKF estimation errors and the theoretical RMSE lower bounds along the trajectories are shown in (c).
@@
The staircase scenario and our experimental setup are shown in Figures 7a and b. Similar to the experiments in the flight arena, we constructed the UWB TDOA localization system with eight UWB anchors in centralized TDOA mode. The anchor positions were selected manually to cover the entire staircase space while maintaining line-of-sight between each other. In the ESKF algorithm, we increased the variance of UWB measurement noise to $\sigma_{ij}^{2}=0.015$ to address the corrupted UWB measurements encountered in this staircase scenario. We use a Leica total station in the tracking mode, which tracks the prism on the UWB handheld device and provides the position measurement at $5$ Hz for ground truth. To quantify the localization accuracy, we manually moved the handheld device at low speed to prevent the total station from losing track and conducted two experiments walking along the staircase. We demonstrate the ground truth trajectory of the first experiment along with the estimation results in Figure 7a. During the experiments, we moved the handheld device upstairs to the second floor and approached the boundary of the anchor constellation, which resulted in challenging anchor-tag geometry [6].
We compute the theoretical lower bound of the localization RMSE along the experimental trajectories through sensor placement analysis and present them together with the ESKF estimation errors over time in the bottom row of Figure 7c. Although the real-world ESKF estimation errors are considerably higher than the theoretical lower bound due to the challenging radio propagation conditions in the staircase, the theoretical values predicted the localization error peaks by analyzing the anchor-tag geometry. The localization performance degrades significantly around $20$ seconds in trajectory #$\#1$ and around $30$ and $55$ seconds in trajectory #$\#2$, which correspond to the time that we approached the boundary of the anchor constellation. With the assistance of sensor placement analysis, we can evaluate the localization performance for regions or trajectories of interest and effectively predict any degradation in positioning accuracy. This information can be valuable for path planning algorithms of mobile robots, for example.
@@ IV-D A Case Study: Localization in a Multi-room Scenario
To demonstrate the entire process of constructing a UWB TDOA localization system, we select a challenging multi-room scenario, which includes a cafeteria and a narrow hallway, as our case study. We selected the cafeteria and the hallway in the University of Toronto Institute for Aerospace Studies (UTIAS) as the experimental space. The multi-room space dimensions were obtained from the building blueprint with the height measured manually. The 3D layout of the multi-room environment, accurately scaled to reflect the real-world dimensions, is visualized in Figure 8a. We modeled the fridge, counter table, and the vending machine as metal obstacles that will lead to severe NLOS scenarios [6] and ignored the movable tables, chairs, and sofas in the cafeteria. As an illustrative example, we aimed to develop a UWB TDOA localization system with the goal of achieving a theoretical localization RMSE lower bound of $20$ centimeters along a pre-defined trajectory from the center of the cafeteria to the end of the hallway. The desired RMSE lower bound value only serves as one example of this case study and can be varied according to different design purposes. We configured the UWB anchors from Bitcraze into decentralized mode [20], which enables scalability in the number of anchors, to cover the entire space. The anchor positions were constrained to be on the boundary of the 3D space for installation purposes.
@@
It is crucial yet challenging to find an optimal anchor placement to ensure reliable localization performance in such complicated environments. Consequently, we employed the optimal sensor placement algorithm [6] briefly introduced in Section III-B to optimize both the required number of anchors and their corresponding positions to meet the accuracy requirement. We selected $22$ sample points to represent the predefined trajectory and optimize the anchor positions for them. In the sensor placement algorithm, we assume the UWB LOS measurements are unbiased with the default $10$ centimeters standard deviation and ignore multi-path effects. The optimization results indicated that $16$ anchors, arranged into $8$ anchor pairs $\Gamma=\{(1,2),(3,4),\cdots,(15,16)\}$, are required to achieve an average RMSE lower bound of $20$ centimeters over the sample points. The optimized anchor positions and the corresponding localization accuracy heatmap, with lower RMSE indicated with darker color, are visualized in Figure 8b. With the optimized anchor placement, each sampled point maintains LOS to several anchor pairs in this challenging environment to ensure reliable localization. As Bitcraze’s UWB anchors communicate and synchronize among all nearby anchors in decentralized mode, UWB tag will receive out-of-sequence TDOA measurements $d_{ij}\notin\Gamma$. Since the out-of-sequence UWB measurements often suffer from degraded anchor geometries, we inflated the variance for $d_{ij}\notin\Gamma$ as $0.025$ in the ESKF algorithm to fuse these measurements. We also increased the Mahalanobis distance parameter of the chi-squared test to $10$ for outlier rejection to avoid rejecting too many measurements in this challenging environment.
After determining the number of anchors and their positions, we deployed the $16$ UWB anchors in the UTIAS cafeteria and hallway according to the sensor placement design. The deployment of the localization system during the experiments is shown in Figure 9a. As explained in Section III-D, we constructed the inertial frame and surveyed the anchor positions within this frame using a Leica total station. The average RMSE between the deployed anchor positions compared to the designed positions is $22$ centimeters, which is partially due to the inconsistency between the actual construction results and the building blueprints. However, this discrepancy is negligible compared to the overall scale of the space, which spans over $25$ meters.
We follow the same procedure as the staircase experiments in Section IV-C and used the Leica total station to provide ground truth. However, due to the challenging geometry at the tight corners around the doorway, the total station lost track at the intersection of the cafeteria and the hallway. Consequently, we had to conduct experiments in the cafeteria and the hallway separately to obtain reliable ground truth position measurements to quantify the localization accuracy.
We conducted three experiments in the cafeteria and two experiments in the hallway following a predefined trajectory. We summarize all the experimental results and demonstrate the estimated trajectories together with the total station ground truth trajectories in Figure 9b. The corresponding RMSE for each experiment and the theoretical RMSE analysis are summarized in Figure 10a. The average localization RMSE in the cafeteria and the narrow hallway are $24$ centimeters and $34$ centimeters, respectively, leading to an overall average of $28$ centimeters in this multi-room environment.
@@
Although the localization accuracies, considering the low-cost IMU and UWB sensors we used in this challenging environment, are commendable and sufficient for pedestrian localization applications, there remains a gap between theoretical analysis and experimental results. The average theoretical localization lower bound derived from sensor placement analysis is $17$ centimeters. This gap is mainly caused by the differences in the UWB sensor noise model between theoretical analysis and real-world experiments. We present the UWB TDOA measurement error histograms among the five experimental trajectories in Figure 10b. The default zero-mean Gaussian noise distribution $n_{ij}\sim\mathcal{N}(0,0.01)$ under LOS conditions is overlaid onto the histograms for comparison. It can be observed that the UWB measurements were severely corrupted compared to the default zero-mean Gaussian error distribution we used in the simulation, which leads to the gap between the real estimated RMSE and the theoretical analysis. In the sensor placement analysis, we assume that the UWB measurements are unbiased under LOS scenarios and neglect multi-path signal propagation. However, the radio propagation environment was more challenging than the simulation setting, with UWB TDOA measurements corrupted by the intrinsic biases of the low-cost DW1000 UWB module [9] and multi-path propagation, especially in the narrow hallway.
@@ V Conclusion
In this article, we presented a UWB TDOA localization system approached from a system-level perspective. We detailed a fully integrated system, from sensor placement optimization, hardware and software architecture, to real-world deployment and system evaluation. Our system provides accurate and robust localization service while maintaining a cost-effective, lightweight, and portable design. We conducted extensive experiments across various real-world environments, demonstrating its accuracy and robustness in multi-agent pedestrian localization, resilience to sensor occlusion, and the ability to evaluate the performance degradation using sensor placement analysis in a challenging staircase. Furthermore, we selected a multi-room environment, including a cafeteria and a narrow hallway, as a case study to exemplify the deployment of our entire UWB TDOA localization system. We started with sensor placement optimization for a pre-defined trajectory, proceeded to the system deployment, and concluded with an evaluation of the real-world localization performance against theoretical analysis. The experimental results show that the proposed UWB TDOA localization system can achieve around $28$ centimeters positioning accuracy in this challenging multi-room environment spanning over $25$ meters, which sheds light on radio-based, scalable, cost-effective, and lightweight indoor localization solutions.

### v2

@@ IV-A Multi-agent Pedestrian Localization
We first evaluate the accuracy and robustness of our UWB TDOA positioning system in multi-agent pedestrian localization through comprehensive real-world experiments. Eight DWM1000 UWB radios are deployed as anchors in centralized TDOA mode with a round-robin topology $\Gamma=\{(8,1),(1,2),\dots,(7,8)\}$. A detailed explanation of centralized and decentralized TDOA modes using Bitcraze’s UWB anchors can be found in [21]. The anchors are placed at the corners of an open two-story indoor flight arena to maximize coverage (Figure 3b). The UWB TDOA noise standard deviation is set to $\sigma_{ij}=0.1$ m, and a chi-squared test with Mahalanobis threshold $5$ is used for outlier rejection. Although the UWB TDOA localization system can provide positioning services to an unlimited number of agents, we demonstrate the multi-agent pedestrian localization performance using five identically designed handheld devices in our experiments.
@@
Two anchor constellations are evaluated. In the first constellation, four anchors are mounted on tripods at a height of $2.65$ m. Five participants freely walked and ran within the anchor convex hull over three trials. We demonstrate one trial of our experiments together with a comparison of the estimated and the ground truth trajectories for each agent in Figure 4. Across the three experimental trials, our localization system demonstrated consistent pedestrian localization performance for the five agents, achieving an average RMSE of $21$ cm with a small standard deviation of $1.56$ cm. In the second anchor constellation, we mounted the four anchors on the room’s ceiling frame, increasing the z-axis separation to $5.5$ m. Three participants carried the handheld devices and conducted three trials of experiments by walking around the same room. With identical parameters applied in the ESKF algorithm, the average RMSE for the three experiments in the second constellation is $16$ cm with a similar standard deviation of $1.63$ cm.
@@
Through these comparative experiments, we observe that changing the sensor placement alone leads to a notable impact on the localization performance, which highlights the importance of UWB anchor placement. We calculate the theoretical lower bounds of localization RMSE in the aforementioned two anchor constellations using sensor placement analysis. The average theoretical RMSE lower bounds in the first and second anchor constellations are $12$ cm and $8$ cm, respectively. We summarize the average experimental RMSE values for multi-agent pedestrian localization and the theoretical RMSE lower bounds in Figure 5 for comparison. The experimental results agree with the theoretical analysis that sensor placements with lower theoretical RMSE bounds generally yield better localization performance.
@@ IV-B Localization under Sensor Occlusion
One distinctive feature of UWB-Inertial localization is that the sensor set can be partially or completely occluded thanks to the obstacle-penetrating ability of UWB measurements. To demonstrate this feature, we put one UWB handheld device into a backpack (see Figure 6) and conducted three trials of experiments in the aforementioned first anchor constellation. Since the device was completely occluded by the backpack throughout the experiments, we created a motion capture tracking object for the backpack to record the ground truth position data approximately. The average localization RMSE together with the theoretical lower bound is summarized in Figure 5.
The average RMSE from the three experiments is $22$ cm, which closely matches the performance of multi-agent pedestrian localization ($21$ cm) within the same constellation. Although the fabric backpack did not pose a severe NLOS challenge for UWB measurements, conventional localization algorithms relying on cameras or LiDARs are unsuitable for this application. This experiment demonstrates that the UWB-inertial localization system has the potential to provide reliable positioning performance even under complete sensor occlusion.
Refer to caption (a) (b) (c)
Figure 7: The staircase scenario and our experimental setup are shown in (a) and (b). The experimental trajectory #1\#1 and the corresponding ground truth trajectory are shown in (a). The ESKF estimation errors and the theoretical RMSE lower bounds along the trajectories are shown in (c).
@@ IV-C Localization in a Staircase Environment
Realistic working conditions, such as building monitoring or multi-room scenarios, often pose geometric challenges and lead to degradation in UWB TDOA localization performance. Therefore, evaluating localization performance through sensor placement analysis is crucial to avoid areas with poor positioning accuracy. To showcase this ability, we conducted experiments in a geometrically challenging staircase environment. We illustrate how sensor placement analysis effectively captures the degradation of localization accuracy induced by the difficult anchor-tag geometry in staircase positioning scenarios.
@@
The staircase scenario and our experimental setup are shown in Figures 7a and b. Eight anchors were deployed in centralized TDOA mode to cover the staircase while maintaining inter-anchor line-of-sight. In the ESKF algorithm, we increased the variance of UWB measurement noise to $\sigma_{ij}^{2}=0.015$ to account for degraded radio conditions. We use a Leica total station in the tracking mode, which tracks the prism on the UWB handheld device and provides the position measurement at $5$ Hz for ground truth. To prevent the total station from losing track of the prism, we conducted two low-speed experiments walking along the staircase to evaluate the localization performance. We demonstrate the ground truth trajectory of the first experiment along with the estimation results in Figure 7a. During the experiments, we moved the handheld device upstairs to the second floor and approached the boundary of the anchor constellation, which resulted in challenging anchor-tag geometry [22].
@@
We compute the theoretical lower bound of the localization RMSE along the experimental trajectories through sensor placement analysis and present them together with the ESKF estimation errors over time in the bottom row of Figure 7c. Although the real-world ESKF estimation errors are considerably higher than the theoretical lower bound due to the challenging radio propagation conditions in the staircase, the theoretical values predicted the localization error peaks by analyzing the anchor-tag geometry. The localization performance degrades significantly around $20$ seconds in trajectory #1\#1 and around $30$ and $55$ seconds in trajectory #2\#2, which correspond to the time that we approached the boundary of the anchor constellation. These results demonstrate that sensor placement analysis can identify regions of degraded localization performance, providing valuable information for path planning algorithms of mobile robots, for example.
@@ IV-D A Case Study: Localization in a Multi-room Scenario
To demonstrate the entire process of constructing a UWB TDOA localization system, we select a challenging multi-room scenario, which includes a cafeteria and a narrow hallway, as our case study. We selected a cafeteria together with a hallway as the experimental space. The multi-room space dimensions were obtained from the building blueprint with the height measured manually. The 3D layout of the multi-room environment, accurately scaled to reflect the real-world dimensions, is visualized in Figure 8a. We modeled the fridge, counter table, and the vending machine as metal obstacles that will lead to severe NLOS scenarios [22] and ignored the movable tables, chairs, and sofas in the cafeteria. As an illustrative example, we aimed to design a UWB TDOA localization system with the goal of achieving a theoretical localization RMSE lower bound of $20$ cm along a pre-defined trajectory from the center of the cafeteria to the end of the hallway. We configured the UWB anchors from Bitcraze into decentralized mode [21], which enables scalability in the number of anchors, to cover the entire space. The anchor positions were constrained to be on the boundary of the 3D space for installation purposes.
@@
We employed the optimal sensor placement algorithm [22] briefly introduced in Section III-B to optimize both the required number of anchors and their corresponding positions to meet the accuracy requirement. A trajectory represented by $22$ sample points are selected and the anchor positions are optimized for them. We assume the UWB LOS measurements are unbiased with the default $10$ cm standard deviation and ignore multi-path effects. The optimization results indicated that $16$ anchors, arranged into $8$ anchor pairs $\Gamma=\{(1,2),(3,4),\cdots,(15,16)\}$, are required to achieve an average RMSE lower bound of $20$ cm over the sample points. The optimized anchor positions and the corresponding localization accuracy heatmap, with lower RMSE indicated with darker color, are visualized in Figure 8b. With the optimized anchor placement, each sampled point maintains LOS to several anchor pairs in this challenging environment to ensure reliable localization. As Bitcraze’s UWB anchors communicate and synchronize among all nearby anchors in decentralized mode, UWB tag will receive out-of-sequence TDOA measurements $d_{ij}\notin\Gamma$. As the out-of-sequence UWB measurements often suffer from degraded anchor geometries, we inflated the variance for $d_{ij}\notin\Gamma$ as $0.025$ in the ESKF algorithm to fuse these measurements. We also increased the Mahalanobis distance parameter of the chi-squared test to $10$ for outlier rejection in this challenging environment.
@@
After determining the number of anchors and their positions, we deployed the $16$ UWB anchors in the cafeteria and the hallway according to the sensor placement design. The deployment of the localization system during the experiments is shown in Figure 9a. As explained in Section III-D, we constructed the inertial frame and surveyed the anchor positions within this frame using a Leica total station. The average RMSE between the deployed anchor positions compared to the designed positions is $22$ cm, which is partially due to the inconsistency between the actual construction results and the building blueprints. However, this discrepancy is negligible compared to the overall scale of the space, which spans over $25$ m. Due to the challenging geometry at the tight corners around the doorway, the total station lost track at the intersection of the cafeteria and the hallway. Consequently, we had to conduct experiments in the cafeteria and the hallway separately to obtain reliable ground truth position measurements from the Leica total station to quantify the localization accuracy.
We conducted three experiments in the cafeteria and two experiments in the hallway following a predefined trajectory and summarize all the experimental results together with the ground truth trajectories in Figure 9b. The corresponding RMSE for each experiment and the theoretical RMSE analysis are summarized in Figure 10a. The average localization RMSE in the cafeteria and the narrow hallway are $24$ cm and $34$ cm, respectively, leading to an overall average of $28$ cm in this multi-room environment.
Although the localization accuracies, considering the low-cost IMU and UWB sensors we used in this challenging environment, are commendable and sufficient for pedestrian localization applications, there remains a gap between theoretical analysis and experimental results. The average theoretical RMSE lower bound from sensor placement analysis is $17$ cm. This discrepancy primarily arises from mismatches between the assumed UWB noise model and real-world measurements. Figure 10b shows histograms of UWB TDOA residuals across five trajectories, overlaid with the assumed zero-mean Gaussian noise model $n_{ij}\sim\mathcal{N}(0,0.01)$ under LOS conditions. In practice, the measurements exhibit significant corruption due to intrinsic DW1000 biases [24] and multipath propagation, particularly in the narrow hallway—effects that are neglected in the theoretical analysis.
@@ V Conclusion
In this article, we presented a system-level UWB TDOA localization framework, spanning sensor placement optimization, hardware and software integration, and real-world deployment. The proposed system provides accurate and robust localization while remaining cost-effective, lightweight, and portable. Extensive experiments across diverse environments demonstrate reliable multi-agent pedestrian localization, robustness to complete sensor occlusion, and the ability of sensor placement analysis to predict performance degradation in a geometrically challenging staircase. A multi-room case study involving a cafeteria and a narrow hallway further illustrates the complete deployment pipeline, from placement optimization to experimental validation. In this environment spanning over $25$ m, the system achieves an average positioning accuracy of approximately $28$ cm, highlighting the potential of scalable, low-cost radio-based localization for indoor applications.